### v1

@@ 5.2 Correlation
As one of the strongest and earliest discovered DIBs, the correlation of DIB $\lambda 5780$ with $\rm E(B-V)$ is the most studied. Our slope of 0.606 Å$\rm\AA\,mag^{-1}$ is close to the slope of 0.64 Å$\rm\AA\,mag^{-1}$ reported by Vos et al. (2011)  [^9] Vos et al. (2011) utilized the data along the specified sight-lines called “$\sigma$-type” to fit.. The slopes of 0.549 Å$\rm\AA\,mag^{-1}$ computed by Herbig (1975)  [^10] The slope was fitted based on the data excluding the sight-lines towards the CYG and SCO-OPH that have skin effect., 0.505 Å$\rm\AA\,mag^{-1}$ fitted by Friedman et al. (2011)  [^11] The slope is actually computed by the inverse of the slope of 1.98, because Friedman et al. (2011) used $\rm E(B-V)=\textit{a}\times EW+\textit{b}$ as their fitting function., 0.525 Å$\rm\AA\,mag^{-1}$ reported in Raimond et al. (2012), and 0.43 Å$\rm\AA\,mag^{-1}$ derived by Lan et al. (2015)  [^12] Lan et al. (2015) employed a power-law function, i.e., $\rm EW=\textit{a}\times E(B-V)^{\gamma}$, as their fitting reference. Nonetheless, for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, the fitted values of $\gamma$ are 1.0, 0.96 and 0.97, which can approximate the linear function. are all smaller than our slope. Some other works have even obtained the slopes lower than those found in the above works, such as the slope of 0.401 Å$\rm\AA\,mag^{-1}$ from Puspitarini et al. (2013) and the slope of 0.419 Å$\rm\AA\,mag^{-1}$ from Vos et al. (2011)  [^13] The slope was fitted based on the data along the specified sight-lines called “$\zeta$-type”.. There are three aspects that can cause the discrepancy of these slopes. One is the measurement error of the DIBs and the extinction. Apart from Lan et al. (2015), the other studies focused on their data within an absolutely narrow range of EW or $\rm E(B-V)$. For instance, the mean EW of the DIB $\lambda 5780$ in Puspitarini et al. (2013) is about 60 mÅ, with the measurement errors exceeding the EWs themselves (see Figure 6 in their work). Similarly, the samples with large EW in Friedman et al. (2011) hardly influence the fitting and significantly deviate from the fitted curve. Consequently, the fitting results are predominantly driven by the samples with $\rm E(B-V)<0.2$ mag (refer to the enlarged region in Figure 4 of their work). Another reason is the different profile definition of the DIB $\lambda 5780$ (see Section 4.2.1 for more details). Lan et al. (2015) separated the broader DIB $\lambda 5778$ from the DIB $\lambda 5780$ and simultaneously fitted them with a double Gaussian profile, whereas the other works, including ours, treated them as a single DIB. Despite the discovery of the DIB $\lambda 5778$ as early as 1975 (Herbig 1975), the subsequent research on the $\lambda 5780$ often overlooked this blended broad DIB. If the contribution of the DIB $\lambda 5778$ by Lan et al. (2015) is attributed to the DIB $\lambda 5780$, the co-added result will align with the slope obtained in this work. The last key factor is the selection of the sight-lines. Vos et al. (2011) elaborated the influence of $\sigma-$ and $\zeta-$type sight-lines on DIB carrier according to the stars from the upper Scorpius. They concluded that the DIB $\lambda 5780$ in the $\sigma-$type sight-line is stronger than that in the $\zeta-$type (Krelowski & Westerlund 1988; Kos & Zwitter 2013), which may be caused by the different ultraviolet (UV) radiation fields, i.e., the external ionized regions for the $\sigma-$type and the UV-shielded could cores for the $\zeta-$type. It is also worth noting that, compared to the DIBs $\lambda 5797$ and $\lambda 6614$, the dispersion of EWs in each extinction bin for the DIB $\lambda 5780$ increases with $\rm E(B-V)$, which further validates the $\sigma-\zeta$ effect on the DIB $\lambda 5780$. Such a $\sigma-\zeta$ effect on our slope is beyond the scope of this work, but our correlations of the DIB $\lambda 5780$ with $\rm E(B-V)$ offer more generalizable conclusions, and circumvent the sample selection effect (the limited number of sight-lines) and the measurements that are sensitive to their own errors. Furthermore, the large number of samples in this work is likely to encompass more $\sigma-$ and $\zeta-$type sight-lines, which will help us to understand the DIB $\lambda 5780$ along the different sight-lines in a more comprehensive way.
@@
For the DIBs $\lambda 5797$ and $\lambda 6614$, their correlations with $\rm E(B-V)$ resembles those of previous works. Our slope of 0.2 Å$\rm\AA\,mag^{-1}$ for the DIB $\lambda 5797$ is consistent with 0.219 Å$\rm\AA\,mag^{-1}$ derived by Herbig (1975), 0.174 Å$\rm\AA\,mag^{-1}$ computed by Friedman et al. (2011) and 0.18 Å$\rm\AA\,mag^{-1}$ fitted by Lan et al. (2015). By similar token, the slopes of 0.262 Å$\rm\AA\,mag^{-1}$ from Herbig (1975), 0.216 Å$\rm\AA\,mag^{-1}$ reported in Friedman et al. (2011) and 0.22 Å$\rm\AA\,mag^{-1}$ documented by Lan et al. (2015) for the DIB $\lambda 6614$ are in good agreement with our slope of 0.274 Å$\rm\AA\,mag^{-1}$. Such consistency of the slopes among different works are unsurprising, as Kos & Zwitter (2013) indicated that the DIBs $\lambda 5797$ and $\lambda 6614$ are unaffected by the $\sigma-\zeta$ effect. Additionally, although the DIB $\lambda 5797$ is influenced by the nearby DIB $\lambda 5795$ (Krelowski & Walker 1987; Friedman et al. 2011), this impact is confined to the blue wing of the DIB $\lambda 5797$ and not as significant as the complete overlap of the DIB $\lambda 5778$ with the broad DIB $\lambda 5780$. As to the DIB $\lambda 6614$, it is not blended with any other DIB (Herbig & Soderblom 1982; Galazutdinov et al. 2008), which makes the correlation with $\rm E(B-V)$ more straightforward. The probable cause of the slight difference in slope still stems from the measurement error of the DIBs and the extinction considering weaker EW per unit extinction of the DIBs $\lambda 5797$ and $\lambda 6614$ compared to the DIB $\lambda 5780$.
@@
We also take the same data used to linear fittings to calculate the Pearson correlation coefficient between EW and $\rm E(B-V)$ for the three DIBs, which are 0.54, 0.51 and 0.64 for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively. None of these DIBs reaches a significant level of correlation with extinction, consistent with the conclusions of many recent studies (Friedman et al. 2011; Fan et al. 2017), which is caused by the variable interstellar environments as mentioned before. Furthermore, the differences of the number of data points that are used to compute correlation coefficients can also lead to the discrepancy of the interpretation of correlation. To generalize our conclusions further, besides the 6,514 common sources, we perform the same fitting procedure on all HQ sources at heliocentric distances greater than 500 pc. The results are also summarized in Table 4, which are comparable to those of the common sources.
5.2.2 DIB V.S. DIB
@@
DIB $\it\lambda 5780-\lambda 5797$: This pair of DIBs has been widely studied in the literature since the first discovery of the DIBs $\lambda 5780$ and $\lambda 5797$ (Heger 1922). The general ratio of $\rm EW_{5797}/EW_{5780}$ is 0.263, which is lower than the value of 0.384 reported by Friedman et al. (2011). The high scatter of the data points used for fitting in Friedman et al. (2011) is the main cause of this discrepancy. In Figure 10, i.e., the figure showing the fitting results of $\rm EW_{5797}$ and $\rm EW_{5780}$, of their work, the fitting is predominantly driven by their samples with $\rm EW_{5780}$ less than 140 mÅ (see the inset of their Figure 10), but approximately one-third of the data points lie below the fitted line. In particular, there is a strong deviation trend for the samples with $\rm EW_{5780}$ between 150 and 300 mÅ. Additionally, our fitting is performed on the medians within each bin, rather than on all data points together as done by Friedman et al. (2011). Results of the later approach is often more susceptible to outliers. The famous $\sigma-\zeta$ effect (see the discussion in Section 5.2.1) can be a probable cause of such a high scatter in Friedman et al. (2011), as seen, for instance, in the sight-line towards Herschel 36 and 6 Cas annotated in their Figure 10. It is also worth noting that while Friedman et al. (2011) used the OB stars within the Galactic plane, our sample includes the sources at higher Galactic latitudes, which can also contribute to the statistical differences between the two studies. The ratio of $\rm EW_{5797}/EW_{5780}$, however, are more complex than hitherto thought, and cannot be adequately represented by a simple linear relationship. In addition to the $\sigma-\zeta$ effect, Farhang et al. (2019) proposed that the carrier of the DIB $\lambda 5780$ can resist X-ray photodissociation and sputtering by fast ions, whereas the carrier of the DIB $\lambda 5797$ breaks down. A more sophisticated method, such as a 3D projection map (Alves et al. 2020; Saydjari et al. 2023), is anticipated to depict it accurately.
@@
DIB $\it\lambda 5780-\lambda 6614$: Among all three DIB pairs, this pair shows the most significant correlation ranked by the Pearson correlation coefficient of 0.788. Interestingly, and somewhat paradoxically, many studies have provided conflicting conclusions regarding their classification into specific families or groups according to the different perspectives. For example, based on their intensity behavior through different types of clouds ($\sigma-$ and $\zeta-$types), Kos & Zwitter (2013) contended that the DIB $\lambda 6614$ belongs to type I (where the linear relations with extinction are similar) and the DIB $\lambda 5780$ is type II (where they changed substantially). In contrast, Fan et al. (2022) categorized the DIBs $\lambda 5780$ and $\lambda 6614$ into the same $\sigma-$DIB group using a hierarchical agglomerative clustering method based on EW of DIB normalized to extinction. Additionally, in the studies conducted by Moutou et al. (1999), Friedman et al. (2011) and Smith et al. (2021), including ours, the correlation of this pair ranked among the top across all DIB pairs. One reasonable explanation for the conflict is that the carriers of the DIBs $\lambda 5780$ and $\lambda 6614$ display different properties in various physical environments, as Fan et al. (2022) argued that the DIBs form a rather continuous sequence. It is worth emulating a multi-dimensional approach by Lan et al. (2015) to decouple different effects on DIB. Through combining extra information such as the sky distribution and kinematics with the ISM tracers, a more exhaustive understanding of DIB will be achieved.
@@
DIB $\it\lambda 6614-\lambda 5797$: There is no evidence that supports the close relationship between the DIBs $\lambda 6614$ and $\lambda 5797$ (Moutou et al. 1999; Friedman et al. 2011; Smith et al. 2021), as the Pearson correlation coefficient is 0.673, ranking the lowest among the three DIB pairs. Although the DIBs $\lambda 5780$ and $\lambda 5797$ unaffected by the $\sigma-\zeta$ effect (Vos et al. 2011), Fan et al. (2022) suggested that they belong to distinct DIB families.
@@
Figure 11: As Figure 10, but for the Gaussian KDE number density plot of EW V.S. EW. From the top panel to the bottom panel, the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are compared with each other. For each panel, the red solid line represents a linear fit to the median EW in each $\rm\Delta EW=0.05\ \AA$ bin. Except for the bottom panel, the bin size of the EW is $\rm\Delta EW=0.02\ \AA$. The coefficients of the linear fit and the Pearson correlation coefficient between the EWs are also annotated in each panel.
@@ 5.3 Central Wavelength in the Rest Frame
The central wavelength of DIB in the rest frame  [^14] In this work, the central wavelengths of the DIBs are all expressed in air and in the heliocentric rest frame without any correction of the RVs of the background stars except for validation of the stellar absorptions in Section 4.2. is one of the most important properties, which can identify the DIB carrier through the match of the band wavelengths and the strength ratios between the observational and laboratory measurements (Campbell et al. 2015; Campbell & Maier 2018; Lallement et al. 2018; Cordiner et al. 2019). Table 5 provides the central wavelength of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ from five canonical DIB catalogs, which are measured using the Doppler shift of the well-identified interstellar atomic or molecular lines, e.g., Na  I D doublet, K  I line and CH line (see Table 2 in Jenniskens & Desert 1994). This traditional method is based on the strong assumption that the DIB carriers co-move with the atoms or molecules so that they share the same Doppler shift. However, all the five catalogs focused on a single or a dozen of sight-lines, and there are probably multiple ISM components along the sight-line, which can lead to the confusion of the ISM Doppler shift (Herbig & Soderblom 1982). Another method to measure the central wavelength of DIB is based on an empirical assumption that, in the Local Standard of Rest (LSR), the matter moving towards the Galactic center (GC) or the Galactic anti-center (GAC) tends to be at rest, meaning the statistical radial velocity of the matter approaches zero (Bovy et al. 2012). Without the interstellar reference lines, Zasowski et al. (2015) were first to successfully measure the central wavelength of the DIB at 1.5273 $\mu$m using this statistical method, so did Zhao et al. (2021a), Gaia Collaboration et al. (2023b) and Saydjari et al. (2023) for the DIB $\lambda 8621$.
@@
Although the spectral wavelength range of LAMOST LRS includes the interstellar absorption lines such as the Ca  II K line at 3933.66 Å and Na  I D doublet at 5889.95 and 5895.92 Å, these lines are prone to blending with stellar absorption lines at the same wavelengths due to the cool stars. Additionally, owing to the increasing modernization around the observation site, the light emitted by high-pressure sodium lamps used for street lighting has contaminated the Na  I D doublet in the spectra (private communication, Chao Liu, 2023). Poznanski et al. (2011) also concluded that Na  I D absorption at low resolution is a bad proxy for extinction. As a consequence, the statistical method instead of the ISM Doppler shift method is employed in this work, but we plan to compare the differences in the central wavelength of DIB between the two methods when the measurements of DIB from hot stars are available in the future. The samples with $|b|<2^{\circ}$, $170^{\circ}<\ell<190^{\circ}$, $d\leq 4\ \rm kpc$, $\rm err(\mu_{mc})<0.5\ \AA$, and $\rm err(RV)<5\ km\,s^{-1}$ are selected to fit the central wavelength of the DIBs. For the DIB $\lambda 5797$, we loosen the constraint of $|b|<2^{\circ}$ to $|b|<10^{\circ}$ because of its limited number of the HQ sources. There are 4934, 1542 and 3186 samples in the GAC for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively. Figure 12 shows the observed central wavelength of DIB as a function of the angular distance from the GAC. By the linear fit to the medians of each bin ($\Delta\ell=1^{\circ}$) using the least squares method, the observed central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ at $\ell=180^{\circ}$ are $5780.68\pm 0.01$, $5797.14\pm 0.02$ and $6613.87\pm 0.01$, respectively. We further take the effect of the solar motion into account, and the final central wavelength $\lambda_{0}$ can be expressed as $\lambda_{\rm obs}\times c/(c+U_{\odot})$, where $\lambda_{\rm obs}$ is the observed central wavelength, $c$ is the speed of light, and $U_{\odot}=10.6\ \rm km\,s^{-1}$ is the radial solar motion (Reid et al. 2019). As listed in Table 5, the final central wavelengths $\lambda_{0}$ of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are $5780.48\pm 0.01$, $5796.94\pm 0.02$ and $6613.64\pm 0.01$, respectively, which are all consistent with the central wavelengths from the cross-reference. Nonetheless, the central wavelength of the DIB $\lambda 5797$ is slightly lower than those from the cross-reference. The obtained slopes are also kind of smaller than the value of $23\pm 3.4\ \rm m\AA\,deg^{-1}$ derived by Gaia Collaboration et al. (2023b) and the value of $19\ \rm m\AA\,deg^{-1}$ estimated by Saydjari et al. (2023) for the DIB $\lambda 8621$ in the GAC, but are quite lower than the value of $57\pm 8\ \rm m\AA\,deg^{-1}$ reported by Zasowski et al. (2015) for the DIB at 1.5273 $\mu$m.
@@
Figure 12: Observed central wavelength ($\lambda_{\rm obs}$) of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ as a function of the angular distance ($\Delta\ell$) from the Galactic anti-center, from top to bottom. The gray dots and their error bars represent the measurements of the central wavelength of DIB and the corresponding uncertainties. The white dots and their black error bars denote the medians and the standard deviations in each $\Delta\ell=1^{\circ}$ bin. The red lines are the linear fits to the medians of each bin using the least squares method. The fitting results, i.e., the slopes $k$, the intercepts $b$, and the final central wavelengths $\lambda_{0}$ considering the solar motion, are annotated in each panel.

### v2

@@ 5.2.1 Correlation between DIBs and E(B-V)
As one of the strongest and earliest discovered DIBs, the correlation of DIB $\lambda 5780$ with $\rm E(B-V)$ is the most studied. Our slope of 0.606 Å$\rm\AA\,mag^{-1}$ is close to the slope of 0.64 Å$\rm\AA\,mag^{-1}$ reported by Vos et al. (2011)  [^9] Vos et al. (2011) utilized the data along the specified sight lines called “$\sigma$-type” to fit.. The slopes of 0.549 Å$\rm\AA\,mag^{-1}$ computed by Herbig (1975)  [^10] The slope was fitted based on the data excluding the sight lines towards the CYG and SCO-OPH that have skin effect., 0.505 Å$\rm\AA\,mag^{-1}$ fitted by Friedman et al. (2011)  [^11] The slope is actually computed by the inverse of the slope of 1.98, because Friedman et al. (2011) used $\rm E(B-V)=\textit{a}\times EW+\textit{b}$ as their fitting function., 0.525 Å$\rm\AA\,mag^{-1}$ reported in Raimond et al. (2012), and 0.43 Å$\rm\AA\,mag^{-1}$ derived by Lan et al. (2015)  [^12] Lan et al. (2015) employed a power-law function, that is, $\rm EW=\textit{a}\times E(B-V)^{\gamma}$, as their fitting reference. Nonetheless, for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, the fitted values of $\gamma$ are 1.0, 0.96, and 0.97, which can approximate the linear function. are all smaller than our slope. Some other works have even obtained the slopes lower than those found in the above works, such as the slope of 0.401 Å$\rm\AA\,mag^{-1}$ from Puspitarini et al. (2013) and the slope of 0.419 Å$\rm\AA\,mag^{-1}$ from Vos et al. (2011)  [^13] The slope was fitted based on the data along the specified sight lines called “$\zeta$-type”.. There are three aspects that can cause the discrepancy of these slopes. One is the measurement error of the DIBs and the extinction. Apart from Lan et al. (2015), the other studies focused on their data within an absolutely narrow range of EW or $\rm E(B-V)$. For instance, the mean EW of the DIB $\lambda 5780$ in Puspitarini et al. (2013) is about 60 mÅ, with the measurement errors exceeding the EWs themselves (see Fig. 6 in their work). Similarly, the samples with large EW in Friedman et al. (2011) hardly influence the fitting and significantly deviate from the fitted curve. Consequently, the fitting results are predominantly driven by the samples with $\rm E(B-V)<0.2$ mag (refer to the enlarged region in Fig. 4 of their work). Another reason is the different profile definition of the DIB $\lambda 5780$ (see Sect. 4.2.1 for more details). Lan et al. (2015) separated the broader DIB $\lambda 5778$ from the DIB $\lambda 5780$ and simultaneously fitted them with a double Gaussian profile, whereas the other works, including ours, treated them as a single DIB. Despite the discovery of the DIB $\lambda 5778$ as early as 1975 (Herbig 1975), the subsequent research on the $\lambda 5780$ often overlooked this blended broad DIB. If the contribution of the DIB $\lambda 5778$ by Lan et al. (2015) is attributed to the DIB $\lambda 5780$, the co-added result will align with the slope obtained in this work. The last key factor is the selection of the sight lines. Vos et al. (2011) elaborated the influence of $\sigma-$ and $\zeta-$type sight lines on DIB carrier according to the stars from the upper Scorpius. They concluded that the DIB $\lambda 5780$ in the $\sigma-$type sight line is stronger than that in the $\zeta-$type (Krelowski & Westerlund 1988; Kos & Zwitter 2013), which may be caused by the different ultraviolet (UV) radiation fields, that is, the external ionized regions for the $\sigma-$type and the UV-shielded could cores for the $\zeta-$type. It is also worth noting that, compared to the DIBs $\lambda 5797$ and $\lambda 6614$, the dispersion of EWs in each extinction bin for the DIB $\lambda 5780$ increases with $\rm E(B-V)$, which further validates the $\sigma-\zeta$ effect on the DIB $\lambda 5780$. Such a $\sigma-\zeta$ effect on our slope is beyond the scope of this work, but our correlations of the DIB $\lambda 5780$ with $\rm E(B-V)$ offer more generalizable conclusions, and circumvent the sample selection effect (the limited number of sight lines) and the measurements that are sensitive to their own errors. Furthermore, the large number of samples in this work is likely to encompass more $\sigma-$ and $\zeta-$type sight lines, which will help us to understand the DIB $\lambda 5780$ along the different sight lines in a more comprehensive way.
@@
For the DIBs $\lambda 5797$ and $\lambda 6614$, their correlations with $\rm E(B-V)$ resembles those of previous works. Our slope of 0.2 Å$\rm\AA\,mag^{-1}$ for the DIB $\lambda 5797$ is consistent with 0.219 Å$\rm\AA\,mag^{-1}$ derived by Herbig (1975), 0.174 Å$\rm\AA\,mag^{-1}$ computed by Friedman et al. (2011), and 0.18 Å$\rm\AA\,mag^{-1}$ fitted by Lan et al. (2015). By similar token, the slopes of 0.262 Å$\rm\AA\,mag^{-1}$ from Herbig (1975), 0.216 Å$\rm\AA\,mag^{-1}$ reported in Friedman et al. (2011) and 0.22 Å$\rm\AA\,mag^{-1}$ documented by Lan et al. (2015) for the DIB $\lambda 6614$ are in good agreement with our slope of 0.274 Å$\rm\AA\,mag^{-1}$. Such consistency of the slopes among different works are unsurprising, as Kos & Zwitter (2013) indicated that the DIBs $\lambda 5797$ and $\lambda 6614$ are unaffected by the $\sigma-\zeta$ effect. Additionally, although the DIB $\lambda 5797$ is influenced by the nearby DIB $\lambda 5795$ (Krelowski & Walker 1987; Friedman et al. 2011), this impact is confined to the blue wing of the DIB $\lambda 5797$ and not as significant as the complete overlap of the DIB $\lambda 5778$ with the broad DIB $\lambda 5780$. As to the DIB $\lambda 6614$, it is not blended with any other DIB (Herbig & Soderblom 1982; Galazutdinov et al. 2008), which makes the correlation with $\rm E(B-V)$ more straightforward. The probable cause of the slight difference in slope still stems from the measurement error of the DIBs and the extinction considering weaker EW per unit extinction of the DIBs $\lambda 5797$ and $\lambda 6614$ compared to the DIB $\lambda 5780$.
@@
We also take the same data used to linear fittings to calculate the Pearson correlation coefficient between EW and $\rm E(B-V)$ for the three DIBs, which are 0.54, 0.51, and 0.64 for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively. None of these DIBs reaches a significant level of correlation with extinction, consistent with the conclusions of many recent studies (Friedman et al. 2011; Fan et al. 2017), which is caused by the variable interstellar environments as mentioned before. Furthermore, the differences of the number of data points that are used to compute correlation coefficients can also lead to the discrepancy of the interpretation of correlation. To generalize our conclusions further, besides the 6,514 common HQ sources, we perform the same fitting procedure on all HQ sources at heliocentric distances greater than 500 pc. The results are also summarized in Table 4, which are comparable to those of the common HQ sources.
5.2.2 Correlations between diffuse interstellar bands
@@
DIB $\lambda 5780-\lambda 5797$: This pair of DIBs has been widely studied in the literature since the first discovery of the DIBs $\lambda 5780$ and $\lambda 5797$ (Heger 1922). The general ratio of $\rm EW_{5797}/EW_{5780}$ is 0.263, which is lower than the value of 0.384 reported by Friedman et al. (2011). The high scatter of the data points used for fitting in Friedman et al. (2011) is the main cause of this discrepancy. In Fig. 10 of Friedman et al. (2011), namely the figure showing the fitting results of $\rm EW_{5797}$ and $\rm EW_{5780}$, the fitting is predominantly driven by their samples with $\rm EW_{5780}$ of less than 140 mÅ (see the inset of their Fig. 10), but approximately one-third of the data points lie below the fitted line. In particular, there is a strong deviation trend for the samples with $\rm EW_{5780}$ between 150 and 300 mÅ. Additionally, our fitting is performed on the medians within each bin, rather than on all data points together as done by Friedman et al. (2011). Results of the later approach is often more susceptible to outliers. The famous $\sigma-\zeta$ effect (see the discussion in Sect. 5.2.1) can be a probable cause of such a high scatter in Friedman et al. (2011), as seen, for instance, in the sight line towards Herschel 36 and 6 Cas annotated in their Fig. 10. It is also worth noting that while Friedman et al. (2011) used the OB stars within the Galactic plane, our sample includes the sources at higher Galactic latitudes, which can also contribute to the statistical differences between the two studies. The ratio of $\rm EW_{5797}/EW_{5780}$, however, are more complex than hitherto thought, and cannot be adequately represented by a simple linear relationship. In addition to the $\sigma-\zeta$ effect, Farhang et al. (2019) proposed that the carrier of the DIB $\lambda 5780$ can resist X-ray photodissociation and sputtering by fast ions, whereas the carrier of the DIB $\lambda 5797$ breaks down. A more sophisticated method, such as a 3D projection map (Alves et al. 2020; Saydjari et al. 2023), is anticipated to depict it accurately.
@@
DIB $\lambda 5780-\lambda 6614$: Among all three DIB pairs, this pair shows the most significant correlation ranked by the Pearson correlation coefficient of 0.788. Interestingly, and somewhat paradoxically, many studies have provided conflicting conclusions regarding their classification into specific families or groups according to the different perspectives. For example, based on their intensity behavior through different types of clouds ($\sigma-$ and $\zeta-$types), Kos & Zwitter (2013) contended that the DIB $\lambda 6614$ belongs to type I (where the linear relations with extinction are similar) and the DIB $\lambda 5780$ is type II (where they changed substantially). In contrast, Fan et al. (2022) categorized the DIBs $\lambda 5780$ and $\lambda 6614$ into the same $\sigma-$DIB group using a hierarchical agglomerative clustering method based on EW of DIB normalized to extinction. Additionally, in the studies conducted by Moutou et al. (1999), Friedman et al. (2011), and Smith et al. (2021), including ours, the correlation of this pair ranked among the top across all DIB pairs. One reasonable explanation for the conflict is that the carriers of the DIBs $\lambda 5780$ and $\lambda 6614$ display different properties in various physical environments, as Fan et al. (2022) argued that the DIBs form a rather continuous sequence. It is worth emulating a multidimensional approach by Lan et al. (2015) to decouple different effects on DIB. Through combining extra information such as the sky distribution and kinematics with the ISM tracers, a more exhaustive understanding of DIB will be achieved.
@@
DIB $\lambda 6614-\lambda 5797$: There is no evidence that supports the close relationship between the DIBs $\lambda 6614$ and $\lambda 5797$ (Moutou et al. 1999; Friedman et al. 2011; Smith et al. 2021), as the Pearson correlation coefficient is 0.673, ranking the lowest among the three DIB pairs. Although the DIBs $\lambda 5780$ and $\lambda 5797$ unaffected by the $\sigma-\zeta$ effect (Vos et al. 2011), Fan et al. (2022) suggested that they belong to distinct DIB families.
@@
Figure 11: As in Fig. 10, but for the Gaussian KDE number density plot of EW V.S. EW. From the top panel to the bottom panel, the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ are compared with each other. For each panel, the red solid line represents a linear fit to the median EW in each $\rm\Delta EW=0.05\ \AA$ bin. Except for the bottom panel, the bin size of the EW is $\rm\Delta EW=0.02\ \AA$. The coefficients of the linear fit and the Pearson correlation coefficient between the EWs are also annotated in each panel.
@@ 5.3 Central wavelength in the rest frame
The central wavelength of DIB in the rest frame  [^14] In this work, the central wavelengths of the DIBs are all expressed in air and in the heliocentric rest frame without any correction of the RVs of the background stars except for validation of the stellar absorptions in Sect. 4.2. is one of the most important properties, which can identify the DIB carrier through the match of the band wavelengths and the strength ratios between the observational and laboratory measurements (Campbell et al. 2015; Campbell & Maier 2018; Lallement et al. 2018; Cordiner et al. 2019). Table 5 provides the central wavelength of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ from five canonical DIB catalogs, which are measured using the Doppler shift of the well-identified interstellar atomic or molecular lines, such as Na  I D doublet, K  I line, and CH line (see Table 2 in Jenniskens & Desert 1994). This traditional method is based on the strong assumption that the DIB carriers co-move with the atoms or molecules so that they share the same Doppler shift. However, all the five catalogs focused on a single or a dozen of sight lines, and there are probably multiple ISM components along the sight line, which can lead to the confusion of the ISM Doppler shift (Herbig & Soderblom 1982). Another method to measure the central wavelength of DIB is based on an empirical assumption that, in the Local Standard of Rest (LSR), the matter moving towards the Galactic center (GC) or the Galactic anti-center (GAC) tends to be at rest, meaning the statistical radial velocity of the matter approaches zero (Bovy et al. 2012). Without the interstellar reference lines, Zasowski et al. (2015) were the first to successfully measure the central wavelength of the DIB at 1.5273 $\mu$m using this statistical method, so did Zhao et al. (2021a), Gaia Collaboration et al. (2023b), and Saydjari et al. (2023) for the DIB $\lambda 8621$.
@@
Although the spectral wavelength range of LAMOST LRS includes the interstellar absorption lines such as the Ca  II K line at 3933.66 Å and Na  I D doublet at 5889.95 and 5895.92 Å, these lines are prone to blending with stellar absorption lines at the same wavelengths due to the cool stars. Additionally, owing to the increasing modernization around the observation site, the light emitted by high-pressure sodium lamps used for street lighting has contaminated the Na  I D doublet in the spectra (private communication, Chao Liu, 2023). Poznanski et al. (2011) also concluded that Na  I D absorption at low resolution is a bad proxy for extinction. As a consequence, the statistical method instead of the ISM Doppler shift method is employed in this work, but we plan to compare the differences in the central wavelength of DIB between the two methods when the measurements of DIB from hot stars are available in the future. The samples with $|b|<2^{\circ}$, $170^{\circ}<\ell<190^{\circ}$, $d\leq 4\ \rm kpc$, $\rm err(\mu_{mc})<0.5\ \AA$, and $\rm err(RV)<5\ km\,s^{-1}$ are selected to fit the central wavelength of the DIBs. For the DIB $\lambda 5797$, we loosen the constraint of $|b|<2^{\circ}$ to $|b|<10^{\circ}$ because of its limited number of the HQ sources. There are 4934, 1542, and 3186 HQ sources in the GAC for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively. Figure 12 shows the observed central wavelength of DIB as a function of the angular distance from the GAC. By the linear fit to the medians of each bin ($\Delta\ell=1^{\circ}$) using the least squares method, the observed central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ at $\ell=180^{\circ}$ are $5780.68\pm 0.01$, $5797.14\pm 0.02$, and $6613.87\pm 0.01$, respectively. We further take the effect of the solar motion into account, and the final central wavelength $\lambda_{0}$ can be expressed as $\lambda_{\rm obs}\times c/(c+U_{\odot})$, where $\lambda_{\rm obs}$ is the observed central wavelength, $c$ is the speed of light, and $U_{\odot}=10.6\ \rm km\,s^{-1}$ is the radial solar motion (Reid et al. 2019). As listed in Table 5, the final central wavelengths $\lambda_{0}$ of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ are $5780.48\pm 0.01$, $5796.94\pm 0.02$, and $6613.64\pm 0.01$, respectively, which are all consistent with the central wavelengths from the cross-reference. Nonetheless, the central wavelength of the DIB $\lambda 5797$ is slightly lower than those from the cross-reference. The obtained slopes are also kind of smaller than the value of $23\pm 3.4\ \rm m\AA\,deg^{-1}$ derived by Gaia Collaboration et al. (2023b) and the value of $19\ \rm m\AA\,deg^{-1}$ estimated by Saydjari et al. (2023) for the DIB $\lambda 8621$ in the GAC, but are quite lower than the value of $57\pm 8\ \rm m\AA\,deg^{-1}$ reported by Zasowski et al. (2015) for the DIB at 1.5273 $\mu$m.
@@
Figure 12: Observed central wavelength ($\lambda_{\rm obs}$) of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ as a function of the angular distance ($\Delta\ell$) from the GAC, from top to bottom. The gray dots and their error bars represent the measurements of the central wavelength of DIB and the corresponding uncertainties. The white dots and their black error bars denote the medians and the standard deviations in each $\Delta\ell=1^{\circ}$ bin. The red lines are the linear fits to the medians of each bin using the least squares method. The fitting results, that is, the slopes $k$, the intercepts $b$, and the final central wavelengths $\lambda_{0}$ considering the solar motion, are annotated in each panel.